### v1

# DiaSynth - Synthetic Dialogue Generation Framework

Abstract
The scarcity of domain-specific dialogue datasets across various domains, from academic topics to everyday conversations, limits the development of dialogue systems for various applications. Existing research is often constrained either by dialogue datasets that are too general or by niche domain dialogue datasets whose scale does not match the required scale for training dialogue systems. To address this gap, we introduce DiaSynth - a synthetic dialogue generation framework capable of generating high-quality, contextually rich dialogues across a wide range of domains. Our approach differs from existing frameworks by dynamically generating dialogues that incorporate simulated personas, subtopics, and diverse conversational characteristics, using a Large Language Model (LLM) with Chain of Thought (CoT) reasoning to create contextually rich, domain-specific dialogues that closely mimic natural human interactions. DiaSynth produces tailored dialogues that emulate realistic conversations. We perform our experiments by generating synthetic data using different LLMs and few-shot examples from DialogSum and SAMSum. The pretrained language models fine-tuned on the synthetic data outperform the base models by 16.47%, while the comparison between models fine-tuned on in-domain data and synthetic data shows that the synthetic data is able to capture 90.48% of the distribution of the in-domain data. The quality of the data generated also scales with the size of LLMs. These results validate DiaSynth’s potential as a robust alternative to traditional data collection methods.
1 Introduction
The development of dialogue systems is an important area of research in the field of natural language processing, due to its wide range of applications ranging from customer service chatbots to virtual personal assistants. The effectiveness of these systems rely heavily on the availability of high-quality, large-scale, domain-specific dialogue datasets. Unfortunately, the lack of high-quality, large-scale datasets across various domains such as academic discussions, healthcare consultations and everyday conversations, poses a big challenge. This scarcity limits the ability of developing dialogue systems that can generalize well across diverse settings and domains.
Existing research has attempted to address this issue by collecting domain specific and real life conversations but they fall into one of the two problematic categories. On the one hand, the conversations in a domain specific dataset are superficial and do not go deep into the domain. On the other hand, niche domain dialogue datasets, while contextually rich, often suffer from limited scale. This imbalance does not allow us to develop dialogue systems for underrepresented domains, where collecting large-scale, high-quality data is either costly or logistically complex.
To address these problems, we introduce DiaSynth, a synthetic dialogue generation framework that produces contextually rich and realistic dialogues tailored to specific domains. DiaSynth, with a Large Language Model (LLM) as the backbone, produces high-quality conversations by simulating personas (representing the characteristics and perspectives of a person) and characteristics, which emulates the implicit qualities of a conversation, such as tone, formality, and agreement levels. By using an LLM as the backbone and Chain of Thoughts (CoT) Wei et al. (2024) as the reasoning mechanism, DiaSynth dynamically generates dialogues that closely mimic real-world conversations, allowing for the rapid generation of tailored dialogue data across a wide range of domains. This approach not only addresses the issue of lack of domain specific dialogue data but also provides a scalable, cost-effective alternative to traditional data collection methods.
To validate the effectiveness of DiaSynth, we evaluated the framework on two criteria: the quality of the generated data and the usability of this data for downstream tasks. The results for the quality criterion showed that the data quality improved with the scale of the model. For the usability criterion, we tested the framework on dialogue summarization task. The results demonstrate that models fine-tuned on DiaSynth-generated data outperformed their base versions by an average of 16.47% across multiple evaluation metrics. Additionally, our comparative analysis reveals that synthetic data captures approximately 90.48% of the performance achieved when fine-tuning on in-domain data, underscoring the framework’s potential as a strong alternate when domain-specific data is unavailable.
The remainder of this paper is organized as follows: Section 2 reviews related work on dialogue datasets and synthetic data generation techniques. Section 3 details the DiaSynth framework, including its architecture and data generation methodology. Section 4 presents our experimental setup, datasets, and evaluation metrics. Section 5 discusses the results of our experiments, comparing the performance of models fine-tuned on DiaSynth-generated data versus in-domain data. Section 6 discusses the limitations of DiaSynth and potential future directions for this research. Finally, Section 7 concludes with a summary of our findings.
2 Related Work
2.1 Personality in Synthetic Data Generation
In recent years, there has been a significant increase in research focused on synthetic dialogue generation, largely driven by advancements in Large Language Models (LLMs). To generate realistic and diverse synthetic data, researchers have incorporated personalities, profiles, and character information when prompting LLMs to generate dialogues Han et al. (2024). By enhancing dialogue realism through the simulation of various personality profiles, utilizing the Big Five personality model, and employing structured prompts, this approach has improved task performance in models fine-tuned on these generated dialogues compared to those trained on general chit-chat datasets.
Moreover, integrating personas into synthetic data generation prompts Chan et al. (2024) has demonstrated that models fine-tuned on personalized synthetic data outperform some LLMs of much larger scales. The inclusion of personas in prompts provides diversity in difficulty levels and ranges within the synthetic data, enabling the models to handle situations of varying complexity.
Our approach involves persona extraction after generating subtopics related to the general topics. This enables the generated dialogues to be more in-depth and specific to those subtopics, enhancing both the scale and quality of domain-specific dialogue generation.
2.2 Prompting Task-Oriented Dialogue Generation
Prompt-based techniques have also emerged as powerful methods for generating high-quality synthetic dialogues, particularly for task-oriented dialogue systems. Steindl et al. (2023) explore the generation of synthetic dialogues from structured prompts, focusing on enhancing task-oriented dialogue systems. Their work demonstrates that prompt engineering can produce dialogues that are contextually appropriate and improve system performance by aligning synthetic data more closely with real-world requirements.
To achieve a higher quantity, diversity, and creativity in human-written instruction data, Wang et al. (2022) propose inputting prompts to LLMs to generate instructions based on a small set of seed human-written instructions. This approach aligns the expanded training data more closely with desired task objectives and allows for iterative improvements, producing more nuanced and effective dialogues that meet specific task demands.
Similarly, our study expands topics into subtopics, ensuring that the generated dialogues provide more in-depth and high-quality conversations. By doing so, we aim to produce synthetic data that not only covers a broader range of scenarios but also delves deeper into each topic, thereby enhancing the overall effectiveness of the dialogue systems trained on this data.
3 DiaSynth
DiaSynth is a synthetic dialogue generation framework designed to address the scarcity of high-quality, large-scale, domain-specific dialogue datasets. The framework functions with an LLM as the backbone and CoT as the reasoning mechanism, to simulate diverse real world dialogues that capture the complexity and nuances of human conversations.
DiaSynth operates by taking in a list of topics from the user based on which the conversations should be generated. The users can optionally provide few-shot examples of the format in which they want the dialogue to be generated. Dialogues that are generated directly from the given list of topics will be too superficial since the topics given by the user will not be specific. To overcome this lack of specificity, we generate $m$ sub topics for each of the $n$ topics given by the user. Generating dialogues from the subtopics will have specificity but the dialogues will lack variety. This is because every dialogue is influenced implicitly by the personas of the people involved in the dialogue and, other characteristics such as the location, emotion and more. To increase both variety and depth in the dialogues generated, we generate $p$ personas for each of the $m$ subtopics. The dialogues are then generated for all possible combinations of personas belonging to each of the subtopics. To further ground the dialogues in various settings and characteristics, we employ CoT reasoning during the generation process. DiaSynth employs CoT to reason about the settings and characteristics of a dialogue, as listed in Table 1, ensuring that the dialogues are contextually rich and realistic. This three-stage pipeline not only guarantees the quality of the generated dialogues but also allows for exponential scalability, as illustrated programmatically in Appendix A.
3.1 Subtopic Generation
Subtopic generation is a crucial step in DiaSynth’s pipeline, since it enhances the specificity and depth of the dialogues that will be generated later. For each primary topic given by the user, DiaSynth generates multiple subtopics, effectively narrowing the focus of the conversation. This breakdown is necessary because the primary topics are often too general to generate contextually rich dialogues on their own. For instance, a topic like “healthcare” can be expanded into subtopics such as “doctor-patient consultations,” “mental health discussions,” and “medical diagnostics,” each of which offers a more focused context for dialogue generation. To achieve this, DiaSynth prompts an LLM to generate the user specified number of subtopics for each primary topic. We also run a similarity check between each of the subtopics and remove subtopics that are too similar to other subtopics using a threshold.
3.2 Persona Generation
Personas of the individuals involved in a conversation are primary influencers in determining how a conversation pans out. Using random personas from persona datasets and prompting the LLM to simulate a dialogue between them about a random topic often leads to superficial dialogues that lack depth and contextual richness. To address this issue, DiaSynth generates a user-specified number of personas for each subtopic, ensuring that the personas are conditioned on the subtopic context. This conditioning prompts the LLM to create personas that are most likely to engage in a meaningful dialogue about the subtopic, such as a medical professional and a patient discussing "medical diagnostics" or a researcher and a student talking about "academic publishing." We also run a similarity check for the personas too. The conditioned persona generation is crucial because it ensures that future dialogues will not only be contextually rich but also exhibit a high level of depth. Each dialogue will be between two personas who have relevant expertise or perspectives on the given subtopic, allowing the conversation to explore nuances that would otherwise be missed in a generic dialogue setting. Notably, during our experiments, the similarity checks indicated no significant overlap among the dialogues, suggesting that the inherent diversity of the dialogues stems from DiaSynth’s multi-stage generation pipeline.
Figure 1: Overview of the Dialogue Synthesis Framework: The diagram shows the step-by-step process of generating dialogues from user-provided topics, involving sub-topic generation, persona generation, dialogue generation, and filtering based on calculated metrics for downstream tasks.
3.3 Dialogue Generation
The final stage in DiaSynth’s pipeline is the generation of dialogues, where all the components—subtopics, personas, and characteristics—converge to create contextually rich and realistic conversations. This step uses an LLM as the backbone and CoT as the reasoning mechanism, allowing the model to simulate dialogues that incorporate various aspects of human interaction. In this stage, for each subtopic, DiaSynth generates dialogues by pairing every possible combination of personas within that subtopic. The generation process also takes into account the predefined dialogue characteristics as listed in Table 1—such as emotional state, formality level, and familiarity between speakers—to guide the conversation flow and style. These characteristics are explicitly defined in the CoT prompt, allowing the LLM to generate dialogues that are both contextually appropriate and realistic.
Characteristic Description
Age and Gender Defines the demographic details of both personas, influencing the style, tone, and cultural references in the dialogue.
Familiarity Level Indicates how well the speakers know each other, which affects the formality, trust, and depth of the conversation.
Emotional States Specifies the emotional condition of the speakers (e.g., angry, happy, sad), impacting the tone and progression of the dialogue.
Formality Level Determines the level of politeness or casualness in the dialogue, which depends on context and familiarity between speakers.
Duration of the Conversation Suggests the intended length of the conversation, which can influence the dialogue’s depth and complexity.
Communication Medium Defines the medium of the conversation (e.g., face-to-face, phone, text), which affects the dialogue style and expressions used.
Topic of the Conversation Provides the main subject around which the dialogue is centered, guiding the content and direction of the conversation.
Location of the Conversation Indicates where the dialogue takes place, adding context that can influence the dialogue’s formality and content.
Agreement or Disagreement Shows whether the speakers agree or disagree on the topic, which drives the dialogue dynamics and interaction style.
Natural Dialogue Features Incorporates elements like fillers (e.g., "umm," "uh"), pauses, and slang, adding authenticity and naturalness to the dialogue.
Table 1: Characteristics of the Dialogue for CoT Prompt
4 Experimental Setup
In this section, we detail the experimental setup used to evaluate the effectiveness of DiaSynth. Our evaluation focuses on two criteria - quality of the dialogues generated and usability of the dialogues generated for a downstream task. Quality of the dialogues is evaluated using metrics such as FED, GPTScore, and G-Eval. The usability of the dialogues generated for a downstream task is evaluated by choosing summarization as the downstream task.
4.1 Quality of the dialogues
To evaluate the quality of the dialogues, we employ the metrics that have been developed for evaluating the quality of text generated by LLMs. The metrics are as follows,
  • FED Mehri and Eskenazi (2020) - FED evaluates the quality of dialogue by appending a positive and a negative utterance to the end of the dialogue and measuring the difference in probabilities assigned to the positive and negative utterances.
  • GPTScore Fu et al. (2024) - GPTScore assesses dialogue quality based on various characteristics such as coherence, diversity, and others. The evaluation involves asking an LLM whether a dialogue meets a specific criterion. The score for each criterion is determined by the probability of the LLM responding affirmatively (’Yes’).
  • G-Eval Liu et al. (2023) - G-Eval evaluates dialogue quality by prompting an LLM to rate the dialogue on a scale of 1 to 3 across multiple criteria. The final score is calculated using a weighted average of each rating, with weights derived from the probability distribution of the LLM’s assigned ratings.
To validate the framework across models and also domains, we generate data using three open source LLMs, one closed source LLM and also use few shot examples from two different dialogue datasets. The four LLMs used are - Phi-3 (phi-3-mini-128k-instruct (3.8B)  [^1] Phi-3 HF link) Abdin et al. (2024), InternLM-2.5 (internlm2_5-7b-chat (7.74B) [^2] InternLM-2.5 HF link) Cai et al. (2024), LLaMA-3 (llama3-8b-instruct (8.03B) [^3] LLaMA-3 HF link) Dubey et al. (2024) and GPT-4o. The two different dialogue datasets that were used are DialogueSum Chen et al. (2021) and SAMSum Gliwa et al. (2019)
4.2 Downstream Task - Summarization
To evaluate the usability of the dialogues generated by DiaSynth, we choose summarization as the downstream task. Summarization is an essential application of dialogue systems, where the goal is to generate concise, and contextually relevant summaries. We use three established evaluation metrics—QAGS, BERTScore, and ROUGE-L—to assess the performance of summarization models fine-tuned on DiaSynth-generated data.
  • QAGS Wang et al. (2020) - QAGS (Question Answering Guided Summarization) evaluates the factual consistency of generated summaries by automatically generating questions based on the summary and then comparing the answers obtained from the summary to those from the source dialogue. In our work, instead of generating questions from the summaries, we generate them from the source dialogues, allowing us to directly compare the consistency and accuracy of summaries generated from different datasets containing the same dialogues.
  • BERTScore Zhang* et al. (2020) - BERTScore measures the semantic similarity between the generated summary and the reference summary using contextual embeddings.
  • ROUGE-L Lin (2004) - ROUGE-L is a widely used evaluation metric that measures the overlap of longest common subsequences (LCS) between the generated summary and the reference summary.
We fine-tune pretrained models that perform well on summarization such as DistilBART, BART Lewis et al. (2020), T5 Raffel et al. (2020) and LED Beltagy et al. (2020), on DiaSynth-generated dialogues and evaluate their performance using the above metrics. We evaluate the effectiveness of DiaSynth in two key aspects: first, by assessing the performance improvement of models fine-tuned on DiaSynth-generated data compared to the pretrained models; and second, by measuring the extent to which DiaSynth-generated data reflects real-world data distribution by comparing the performance of models fine-tuned on DiaSynth data versus those fine-tuned on in-domain data.
5 Results
This section discusses the results of the data generated using DiaSynth (quality of the data and usability in downstream tasks) with different LLMs as the backbone and varying few-shot examples. Specifically, we utilized Phi-3, InternLM-2.5, LLaMA-3 and GPT-4o as the LLM backbones, and the few-shot examples were sourced from DialogueSum and SAMSum datasets. These combinations allow us to evaluate the robustness and adaptability of DiaSynth across different models and few shot examples. In total, eight distinct datasets were generated using DiaSynth by pairing each LLM backbone with the two sets of few-shot examples, resulting in all possible combinations. For each dataset, DiaSynth was provided with the same 16 broad topics and tasked with generating 6 subtopics for each topic, followed by creating 6 personas for each subtopic. The statistics of the datasets generated using DiaSynth, including the number of dialogues, average number of turns, and average number of tokens per turn, are summarized in Table 2.
LLM Few-Shot Number of Samples Avg. number of turns Avg. number of tokens per turn Diversity (rouge-l)
Phi-3 DialogueSum 1215 9.13 20.38 0.27
Phi-3 SAMSum 1410 13.98 13.94 0.27
InternLM-2.5 DialogueSum 1035 9.23 27.98 0.30
InternLM-2.5 SAMSum 1135 13.96 19.07 0.29
LLaMA-3 DialogueSum 1154 6.86 31.99 0.29
LLaMA-3 SAMSum 1195 10.54 20.41 0.29
GPT-4o DialogSum 1375 15.16 15.96 0.29
GPT-4o SAMSum 1380 15.43 13.53 to 0.28
Table 2: Data Statistics
5.1 Quality of the Dialogues
The quality of the synthetic datasets produced by DiaSynth was evaluated using FED, GPTScore, and G-Eval metrics, as detailed in Tables 3, 4, and 5. The results reveal distinct variations in performance across different model and dataset configurations, reflecting the unique characteristics of each.
The FED scores in Table 3 indicate that LLaMA-3 and GPT-4o consistently outperform Phi-3 and InternLM-2.5, particularly in coherence, consistency, and diversity, with LLaMA-3 leading in the DialogSum dataset (e.g., coherence 0.9684, consistency 0.9570, diversity 0.9596) and GPT-4o excelling in the SAMSum dataset, especially in coherence (0.9939) and consistency (0.9878). Table 4 further supports LLaMA-3’s adaptability and error-handling abilities, as it outperforms others in diversity (0.0430) and inquisitiveness (0.0415) in DialogSum, while maintaining an edge in SAMSum for diversity (0.0655) and inquisitiveness (0.0622). Though GPT-4o remains strong in coherence and consistency, Phi-3 shows competitive coherence in SAMSum (0.0325). Finally, Table 5 highlights GPT-4o’s dominance in engagingness and naturalness with perfect scores (3.0) for DialogSum, while InternLM-2.5 stands out in coherence (2.9990) and groundedness (2.9973) for DialogSum, and coherence (2.9983) and groundedness (2.9952) for SAMSum, suggesting it maintains high factual accuracy. The contrasting performance of GPT-4o on the DialogSum and SAMSum datasets can be attributed to the differing structures of the dialogues in these datasets. DialogSum consists of more formal and structured dialogues, which aligns with the typical response style of GPT-4o, leading to its stronger performance. In contrast, SAMSum contains more casual, human-like conversations, which might explain GPT-4o’s relatively poorer performance, as it may not adapt as well to the informal, spontaneous nature of such dialogues. Overall, while GPT-4o excels in natural and engaging dialogue, LLaMA-3 offers the most versatility, and InternLM-2.5 provides a strong alternative with high coherence and groundedness.
The observed superiority of LLaMA-3 over GPT-4o is surprising because an 8 billion 8-bit quantized model not only competes with but also performs better than GPT-4o in certain metrics. We hypothesize that this could be due to the way GPT-4o was trained, which might make it more constrained in its responses, whereas LLaMA-3, being an open-source model, operates with fewer restrictions. This allows LLaMA-3 to exhibit greater flexibility, diversity, and adaptability in generating dialogues, potentially explaining its better performance in certain metrics. These characteristics can be seen in criteria like ’inquisitiveness’ and ’likeability’ in Table 4 and, ’depth’ and ’diverse’ in Table 3. These results suggest that for building human-like data generation frameworks, open-source LLMs are a more suitable choice than closed-source LLMs. The minimal constraints on response formatting during the training of open-source models enable them to generate more diverse, flexible, and human-like dialogues, making them better suited for tasks requiring natural and conversational interactions.
coherent error recovery consistent diverse depth likeable understand flexible informative inquisitive
DIALOGUESUM
Phi-3 0.9536 0.9440 0.9540 0.9534 0.9521 -0.0005 0.9353 -3.96E-05 0.0009 -0.0033
InternLM-2.5 0.8439 0.8313 0.8359 0.8353 0.8352 0.0048 0.8278 -0.0046 0.0042 0.0069
LLaMA-3 0.9684 0.9522 0.9570 0.9596 0.9592 0.0032 0.9453 -0.0063 0.0063 0.0105
GPT-4o 0.9525 0.9407 0.9417 0.9423 0.9425 0.0121 0.9368 -0.0027 0.0085 0.0144
SAMSUM
Phi-3 0.9161 0.9088 0.9199 0.9161 0.9130 -0.0004 0.9014 -0.0024 0.0034 -0.0040
InternLM-2.5 0.8746 0.8647 0.8734 0.8655 0.8661 0.0033 0.8582 -0.0019 0.0106 0.0028
LLaMA-3 0.9829 0.9677 0.9757 0.9712 0.9731 0.0003 0.9593 -0.0048 0.0100 0.0029
GPT-4o 0.9939 0.9876 0.9878 0.9836 0.9824 0.0083 0.9788 0.0004 0.0141 0.006
Table 3: FED scores
coherence diversity flexibility understandability inquisitiveness consistency informativeness likeability depth error recovery
DIALOGUESUM
Phi-3 0.0286 0.0310 0.0218 0.0193 0.0363 0.0369 0.0172 0.0213 0.0117 0.0342
InternLM-2.5 0.0069 0.0196 0.0084 0.0061 0.0244 0.0137 0.0148 0.0050 0.0080 0.0197
LLaMA-3 0.0189 0.0430 0.0186 0.0220 0.0415 0.0321 0.0318 0.0110 0.0201 0.0440
GPT-4o 0.0039 0.0156 0.0059 0.0039 0.018 0.0080 0.0097 0.0026 0.0053 0.0135
SAMSUM
Phi-3 0.0325 0.0372 0.0260 0.0270 0.0395 0.0415 0.0201 0.0232 0.0126 0.0290
InternLM-2.5 0.0128 0.0408 0.0194 0.0174 0.0504 0.0306 0.0328 0.0142 0.0177 0.0407
LLaMA-3 0.0288 0.0655 0.0306 0.0365 0.0622 0.0612 0.0542 0.0168 0.0270 0.0558
GPT-4o 0.0055 0.0162 0.0094 0.0084 0.0186 0.0119 0.0129 0.0038 0.0079 0.0199
Table 4: GPTScore
engagingness naturalness coherence groundedness
DIALOGUESUM
Phi-3 2.5236 2.7238 2.6308 2.5557
InternLM-2.5 2.9995 2.9989 2.9990 2.9973
LLaMA-3 2.9987 2.9988 2.9972 2.9935
GPT-4o 3 3 3 2.9975
SAMSUM
Phi-3 2.4623 2.6821 2.5848 2.5060
InternLM-2.5 2.9992 2.9969 2.9983 2.9952
LLaMA-3 2.9976 2.9971 2.9969 2.9916
GPT-4o 2.9994 2.9977 2.9982 2.9944
Table 5: G_EVAL
5.2 Fine-tuning and Performance Results
To validate the usability of the synthetic data generated using DiaSynth, we fine-tuned and evaluated several pretrained language models on the task of dialogue summarization. The summaries for dialogues generated by different LLMs were created using the corresponding LLMs through prompting. The pretrained models used for evaluation include DistilBART, BART, T5, and LED.
The experimental setup is designed as follows:
  • All reported metrics are based on the validation and test sets of DialogSum and SAMSum to maintain consistency in evaluation.
  • To assess the effectiveness of DiaSynth-generated data, we compare the performance of models fine-tuned on DiaSynth data with their corresponding base models, which are evaluated without any fine-tuning.
  • Given that the training sets of DialogSum and SAMSum are larger than the synthetic datasets generated in this study, we randomly sampled instances from the in-domain training sets to match the size of the DiaSynth-generated data. This ensures a fair comparison in terms of training scale.
  • The primary goal of the experiment is to quantify the percentage improvement offered by models fine-tuned on DiaSynth-generated data compared to their base versions and to evaluate the extent to which the generated data capture the distributional characteristics of the in-domain data.
The results presented in Table 6 highlight the performance of each model before and after fine-tuning. Notably, base models such as DistilBART and LED exhibit relatively high QAGS scores before fine-tuning, primarily because these models tend to repetitively generate content similar to the input dialogues, thereby artificially inflating the factual consistency metric. This repetition, while beneficial for preserving factual alignment, does not necessarily translate into improved quality or informativeness of the summaries.
After fine-tuning on DiaSynth-generated data, models show a significant increase in BERTScore and ROUGE-L compared to their base versions, demonstrating DiaSynth’s effectiveness in enhancing model performance for downstream tasks. This result suggests that DiaSynth is valuable for generating dialogue data in domains with limited training resources. Notably, models fine-tuned on GPT-4o-generated data performed best for DialogSum, while those fine-tuned on LLaMA-3-generated data excelled on SAMSum, further reinforcing that open-source LLMs are more adept at generating informal, casual dialogues like those in SAMSum. The percentage improvement of models fine-tuned on LLaMA-3-generated data was calculated using Equation 1, showing an average improvement of 16.47% across the 24 reported results. We chose to calculate on LLaMA-3 because across both quality and usability it performs better than GPT-4o.
Percentage Improvement=Score After Fine-tuningScore Before Fine-tuningScore Before $\text{Percentage Improvement}=\frac{\text{Score After Fine-tuning}-\text{ScoreBefore Fine-tuning}}{\text{Score Before Fine-tuning}}\times 100$ (1)
Coverage Percentage=Finetuned Score on DiaSynth DataFinetuned Score on In-Domain Base $\text{Coverage Percentage}=\frac{\text{Finetuned Score on DiaSynth Data}}{\text{Finetuned Score on In-Domain Base Data}}\times 100$ (2)
To assess the percentage coverage of the distributional characteristics of the in-domain data by the synthetically generated data, we use Equation 2 and the scores of models fine-tuned on LLaMA-3 generated data. Across the 24 reported results, the overall coverage percentage of the LLaMA-3 generated data is 90.48%. Notably, the QAGS scores of models fine-tuned on synthetic data surpass those of models trained on in-domain data, suggesting that synthetic data can match or even exceed in-domain data performance in some aspects. Excluding QAGS, the coverage percentage is calculated to be 77.07%, highlighting that even with moderate LLMs of small scale (3B - 8B), high-quality synthetic dialogue datasets can be effectively created.
Model Before Finetuning Finetuning on Base Data
QAGS BERTScore ROUGE-L QAGS BERTScore ROUGE-L
DIALOGSUM
distillbart-cnn 0.6134 0.5093 0.1950 0.5586 0.7005 0.3367
bart-base 0.7007 0.5274 0.1375 0.4789 0.6868 0.2969
t5-base 0.5901 0.5491 0.1812 0.4766 0.6953 0.2986
led-base-16384 0.8261 0.5471 0.1634 0.4872 0.7084 0.3165
SAMSUM
distillbart-cnn 0.6627 0.5500 0.2394 0.6041 0.6849 0.3578
bart-base 0.7563 0.4389 0.1765 0.5302 0.6520 0.3049
t5-base 0.5574 0.4190 0.1237 0.5460 0.6448 0.3000
led-base-16384 0.7429 0.4310 0.1812 0.5440 0.6522 0.3175
Table 6: Performance of models before and after finetuning on in-domain data
Model Phi-3 InternLM-2.5 LLaMA-3 GPT-4o
QAGS BERTScore ROUGE-L QAGS BERTScore ROUGE-L QAGS BERTScore ROUGE-L QAGS BERTScore ROUGE-L
DIALOGUESUM
distillbart-cnn 0.6588 0.5778 0.2187 0.6420 0.6008 0.2167 0.6586 0.6161 0.2040 0.6713 0.6242 0.2014
bart-base 0.5355 0.5958 0.2029 0.5418 0.6212 0.1897 0.5825 0.6033 0.1789 0.5590 0.6039 0.1769
t5-base 0.5937 0.5949 0.2047 0.5825 0.5941 0.1878 0.6034 0.6172 0.1959 0.6305 0.6319 0.2044
led-base-16384 0.5358 0.6129 0.2109 0.5189 0.6027 0.1606 0.5697 0.6302 0.1999 0.5791 0.6308 0.1989
SAMSUM
distillbart-cnn 0.6585 0.5931 0.2262 0.6388 0.6066 0.2422 0.6849 0.6029 0.2374 0.6757 0.6029 0.2291
bart-base 0.5648 0.5665 0.2146 0.5435 0.5663 0.2021 0.6132 0.5899 0.2345 0.5707 0.5808 0.2154
t5-base 0.5905 0.5397 0.2085 0.5457 0.5193 0.1854 0.6412 0.5054 0.1976 0.6023 0.5419 0.1979
led-base-16384 0.5883 0.5477 0.2289 0.5457 0.5615 0.2167 0.5917 0.5785 0.2390 0.5738 0.569 0.2298
Table 7: Performance metrics
6 Limitations
Despite the promising results, our approach has some limitations. Firstly, different LLMs exhibit varied performance based on the dialogue structure, with certain models like LLaMA-3 performing better for more informal dialogues (e.g., SAMSum) and others like GPT-4o excelling in structured, formal dialogues (e.g., DialogSum). This indicates that there is no single model that can universally handle all types of dialogue structures, requiring a tailored choice of LLM depending on the target application.
Secondly, the generation process may suffer from a lack of knowledge on certain topics, especially in cases where the LLMs were not sufficiently trained on those domains. Additionally, our framework relies on zero-shot generation for personas and sub-topics, which, while flexible, can sometimes result in less coherent or less accurate persona simulation, as it is not fine-tuned for specific contexts. These limitations suggest directions for future work, such as combining LLMs to leverage their strengths or incorporating more topic-specific training to enhance knowledge coverage.
7 Conclusion
In this paper, we introduced DiaSynth, a synthetic dialogue generation framework capable of producing high-quality, contextually rich dialogues across a wide range of domains. Our experiments demonstrated that models fine-tuned on DiaSynth-generated data exhibit significant improvements in downstream tasks, as evidenced by substantial increases in BERTScore and ROUGE-L compared to their base models. These results highlight the potential of DiaSynth as an effective tool for generating dialogue data, particularly in domains where training data is scarce.
Furthermore, our analysis showed that different LLMs excel in different dialogue structures, with LLaMA-3 performing better for informal dialogues and GPT-4o for more structured settings. This insight suggests that leveraging open-source LLMs may be more advantageous for generating human-like conversational data. Despite certain limitations, such as varying LLM performance across dialogue types and knowledge gaps in zero-shot generation, DiaSynth presents a promising approach to dialogue data generation and offers a valuable resource for future advancements in building more sophisticated and adaptable dialogue systems.
Appendix A Scalability of DiaSynth
This section illustrates the scalability of DiaSynth with a python program and examples.
1def calculate_total_dialogs(n, m, p):
2 """
3 Calculate the total number of dialogs generated.
4
5 Parameters:
6 n (int): Number of topics
7 m (int): Number of subtopics per topic
8 p (int): Number of personas per subtopic
9
10 Returns:
11 int: Total number of dialogs generated
12 """
13 # Calculate the total number of subtopics
14 total_subtopics = n * m
15
16 # Calculate the number of dialogs for each subtopic using combinations of personas
17 dialogs_per_subtopic = (p * (p - 1)) // 2
18
19 # Total dialogs generated across all subtopics
20 total_dialogs = total_subtopics * dialogs_per_subtopic
21
22 return total_dialogs
Listing 1: Calculation of Total Dialogs Generated
A.1 Example 1
  • Number of topics ($n$): 10
  • Number of subtopics per topic ($m$): 5
  • Number of personas per subtopic ($p$): 3
Calculation:
Total subtopics $=n\times m=10\times 5=50$
Dialogues per subtopic $=\frac{p\times(p-1)}{2}=\frac{3\times 2}{2}=3$
Total dialogues $=50\times 3=150$
This setup generates 150 dialogues.
A.2 Example 2
  • Number of topics ($n$): 20
  • Number of subtopics per topic ($m$): 4
  • Number of personas per subtopic ($p$): 5
Calculation:
Total subtopics $=20\times 4=80$
Dialogues per subtopic $=\frac{5\times 4}{2}=10$
Total dialogues $=80\times 10=800$
This setup generates 800 dialogues.
A.3 Example 3
  • Number of topics ($n$): 15
  • Number of subtopics per topic ($m$): 6
  • Number of personas per subtopic ($p$): 10
Calculation:
Total subtopics $=15\times 6=90$
Dialogues per subtopic $=\frac{10\times 9}{2}=45$
Total dialogues $=90\times 45=4050$
This setup generates 4050 dialogues.
A.4 Scaling Observations
  • Linear Scaling with Topics and Subtopics: Increasing the number of topics or subtopics results in a linear increase in the total number of dialogues, making it straightforward to expand the scope of dialogue generation.
  • Exponential Scaling with Personas: The number of dialogues scales exponentially as the number of personas increases because each additional persona allows for more combinations, making the framework highly scalable for complex scenarios.
  • Practical Use Case: For specific domains like academic, healthcare, or business, these parameters can be adjusted to generate thousands of dialogues to fit the needs of various applications such as training chatbots, virtual assistants, or dialogue-based assessments.
These examples illustrate DiaSynth’s potential for rapid and scalable generation of dialogues, which can be tailored to different domains by simply adjusting the input parameters.

### v2

@@
-DiaSynth - Synthetic Dialogue Generation Framework
+DiaSynth: Synthetic Dialogue Generation Framework for Low Resource Dialogue Applications
 Abstract
-The scarcity of domain-specific dialogue datasets across various domains, from academic topics to everyday conversations, limits the development of dialogue systems for various applications.
-Existing research is often constrained either by dialogue datasets that are too general or by niche domain dialogue datasets whose scale does not match the required scale for training dialogue systems.
+The scarcity of domain-specific dialogue datasets limits the development of dialogue systems across applications.
+Existing research is constrained by general or niche datasets that lack sufficient scale for training dialogue systems.
 To address this gap, we introduce DiaSynth - a synthetic dialogue generation framework capable of generating high-quality, contextually rich dialogues across a wide range of domains.
-Our approach differs from existing frameworks by dynamically generating dialogues that incorporate simulated personas, subtopics, and diverse conversational characteristics, using a Large Language Model (LLM) with Chain of Thought (CoT) reasoning to create contextually rich, domain-specific dialogues that closely mimic natural human interactions.
-DiaSynth produces tailored dialogues that emulate realistic conversations.
+Unlike existing frameworks, DiaSynth uses Large Language Models (LLMs) and Chain of Thought (CoT) reasoning to generate dynamic, domain-specific dialogues with simulated personas and diverse conversational features.
 We perform our experiments by generating synthetic data using different LLMs and few-shot examples from DialogSum and SAMSum.
-The pretrained language models fine-tuned on the synthetic data outperform the base models by 16.47%, while the comparison between models fine-tuned on in-domain data and synthetic data shows that the synthetic data is able to capture 90.48% of the distribution of the in-domain data.
-The quality of the data generated also scales with the size of LLMs.
+The pretrained language models fine-tuned on the synthetic data outperform the base models by 16.47% on dialogue summarization, while the comparison between models fine-tuned on in-domain data and synthetic data shows that the synthetic data is able to capture 90.48% of the performance distribution of the in-domain data on dialogue summarization.
+The quality of the data generated also increases as we increase the size of LLM from 3B to 8B.
 These results validate DiaSynth’s potential as a robust alternative to traditional data collection methods.
+We open source the code and data generated for future research.
 1 Introduction
-The development of dialogue systems is an important area of research in the field of natural language processing, due to its wide range of applications ranging from customer service chatbots to virtual personal assistants.
-The effectiveness of these systems rely heavily on the availability of high-quality, large-scale, domain-specific dialogue datasets.
-Unfortunately, the lack of high-quality, large-scale datasets across various domains such as academic discussions, healthcare consultations and everyday conversations, poses a big challenge.
-This scarcity limits the ability of developing dialogue systems that can generalize well across diverse settings and domains.
-Existing research has attempted to address this issue by collecting domain specific and real life conversations but they fall into one of the two problematic categories.
+Figure 1: Overview of DiaSynth
+Dialogue systems are crucial in natural language processing due to applications like customer service chatbots and virtual assistants.
+Their effectiveness depends on large, high-quality, domain-specific datasets.
+The lack of large-scale, high-quality datasets across domains like academic discussions, healthcare, and everyday conversations poses a challenge.
+This scarcity limits the development of dialogue systems that generalize well across domains.
+Prior work (Feng et al.
+(2020), Zeng et al.
+(2020), Budzianowski et al.
+(2018)) collects domain-specific dialogues but often lacks depth, scale, or domain diversity.
 On the one hand, the conversations in a domain specific dataset are superficial and do not go deep into the domain.
 On the other hand, niche domain dialogue datasets, while contextually rich, often suffer from limited scale.
-This imbalance does not allow us to develop dialogue systems for underrepresented domains, where collecting large-scale, high-quality data is either costly or logistically complex.
+This imbalance hinders dialogue system development in underrepresented domains, where data collection is costly and complex.
 To address these problems, we introduce DiaSynth, a synthetic dialogue generation framework that produces contextually rich and realistic dialogues tailored to specific domains.
-DiaSynth, with a Large Language Model (LLM) as the backbone, produces high-quality conversations by simulating personas (representing the characteristics and perspectives of a person) and characteristics, which emulates the implicit qualities of a conversation, such as tone, formality, and agreement levels.
-By using an LLM as the backbone and Chain of Thoughts (CoT) Wei et al.
-(2024) as the reasoning mechanism, DiaSynth dynamically generates dialogues that closely mimic real-world conversations, allowing for the rapid generation of tailored dialogue data across a wide range of domains.
-This approach not only addresses the issue of lack of domain specific dialogue data but also provides a scalable, cost-effective alternative to traditional data collection methods.
+DiaSynth, using a Large Language Model (LLM), generates high-quality conversations by simulating personas and conversation characteristics like tone and formality.
+With LLMs and Chain of Thoughts (CoT) Wei et al.
+(2024), DiaSynth generates dialogues that mimic real-world conversations for a wide range of domains.
+CoT plays a crucial part by building different environments for different personas (refer to Appendix D for examples) which influence varied conversations.
+This approach addresses data scarcity and offers a scalable, cost-effective alternative to traditional methods.
 To validate the effectiveness of DiaSynth, we evaluated the framework on two criteria: the quality of the generated data and the usability of this data for downstream tasks.
 The results for the quality criterion showed that the data quality improved with the scale of the model.
-For the usability criterion, we tested the framework on dialogue summarization task.
-The results demonstrate that models fine-tuned on DiaSynth-generated data outperformed their base versions by an average of 16.47% across multiple evaluation metrics.
-Additionally, our comparative analysis reveals that synthetic data captures approximately 90.48% of the performance achieved when fine-tuning on in-domain data, underscoring the framework’s potential as a strong alternate when domain-specific data is unavailable.
-The remainder of this paper is organized as follows: Section 2 reviews related work on dialogue datasets and synthetic data generation techniques.
-Section 3 details the DiaSynth framework, including its architecture and data generation methodology.
-Section 4 presents our experimental setup, datasets, and evaluation metrics.
-Section 5 discusses the results of our experiments, comparing the performance of models fine-tuned on DiaSynth-generated data versus in-domain data.
-Section 6 discusses the limitations of DiaSynth and potential future directions for this research.
-Finally, Section 7 concludes with a summary of our findings.
+For usability, we tested DiaSynth on dialogue summarization.
+Models fine-tuned on DiaSynth data outperformed base versions by 16.47% on average.
+Additionally, DiaSynth’s synthetic data captures 90.48% of in-domain performance, highlighting its potential as a strong alternative when domain-specific data is unavailable.
+The remainder of this paper is organized as follows: Section 2 reviews related work on dialogue datasets and synthetic data.
+Section 3 details the DiaSynth framework and methodology.
+Section 4 describes our experimental setup and evaluation.
+Section 5 compares the performance of models fine-tuned on DiaSynth data to in-domain data.
+Section 6 concludes with a summary of our findings and Section 7 discusses the limitations of DiaSynth and potential future directions for this research.
 2 Related Work
@@
 2.2 Prompting Task-Oriented Dialogue Generation
+LLM Few-shot examples Number of Samples Avg. number of turns Avg. number of tokens per turn Diversity (ROUGE-L)
+Phi-3 DialogSum 1215 9.13 20.38 0.27
+InternLM-2.5 DialogSum 1035 9.23 27.98 0.30
+LLaMA-3 DialogSum 1154 6.86 31.99 0.29
+GPT-4o DialogSum 1375 15.16 15.96 0.29
+Phi-3 SAMSum 1410 13.98 13.94 0.27
+InternLM-2.5 SAMSum 1135 13.96 19.07 0.29
+LLaMA-3 SAMSum 1195 10.54 20.41 0.29
+GPT-4o SAMSum 1380 15.43 13.53 0.28
+Table 1: Data Statistics
 Prompt-based techniques have also emerged as powerful methods for generating high-quality synthetic dialogues, particularly for task-oriented dialogue systems.
@@
 By doing so, we aim to produce synthetic data that not only covers a broader range of scenarios but also delves deeper into each topic, thereby enhancing the overall effectiveness of the dialogue systems trained on this data.
+2.3 Existing Task-Oriented Dialogue Datasets
+In addition to prompt-based synthetic data generation, various large-scale dialogue datasets have been instrumental in advancing task-oriented dialogue systems.
+Among these, the MultiWOZ dataset Budzianowski et al.
+(2018) is a prominent resource, providing richly annotated dialogues across multiple domains.
+MultiWOZ has enabled researchers to train models capable of handling complex, multi-turn interactions across diverse tasks.
+The nature of MultiWOZ’s annotations has made it a benchmark for evaluating the performance of dialogue systems, though it is often complemented by synthetic data to introduce further diversity and variation in dialogue scenarios.
+Similarly, Doc2Dial Feng et al.
+(2020) is another widely used dataset designed specifically for document-grounded dialogue systems.
+Doc2Dial includes conversations grounded in structured documents, focusing on providing users with accurate and relevant information based on their inquiries.
+This dataset has been instrumental in improving the ability of dialogue systems to retrieve and generate accurate responses when interacting with complex information sources.
+However, much like MultiWOZ, Doc2Dial’s scope is limited to the predefined topics and domains covered within the dataset, which can restrict model generalizability to new or unseen situations.
+To overcome the domain-specific limitations of datasets, our study adopts a synthetic data generation approach that expands existing topics into subtopics, thus providing a broader and deeper pool of conversational data.
+By incorporating both task-oriented prompts and personas, our generated dialogues aim to complement these datasets by offering more personalized and contextually rich conversations, thereby enhancing the robustness and versatility of dialogue systems
 3 DiaSynth
+coherent error recovery consistent diverse depth likeable understand flexible informative inquisitive
+DIALOGUESUM
+Phi-3 0.9536 0.9440 0.9540 0.9534 0.9521 -0.0005 0.9353 -3.96E-05 0.0009 -0.0033
+InternLM-2.5 0.8439 0.8313 0.8359 0.8353 0.8352 0.0048 0.8278 -0.0046 0.0042 0.0069
+LLaMA-3 0.9684 0.9522 0.9570 0.9596 0.9592 0.0032 0.9453 -0.0063 0.0063 0.0105
+GPT-4o 0.9525 0.9407 0.9417 0.9423 0.9425 0.0121 0.9368 -0.0027 0.0085 0.0144
+SAMSUM
+Phi-3 0.9161 0.9088 0.9199 0.9161 0.9130 -0.0004 0.9014 -0.0024 0.0034 -0.0040
+InternLM-2.5 0.8746 0.8647 0.8734 0.8655 0.8661 0.0033 0.8582 -0.0019 0.0106 0.0028
+LLaMA-3 0.9829 0.9677 0.9757 0.9712 0.9731 0.0003 0.9593 -0.0048 0.0100 0.0029
+GPT-4o 0.9939 0.9876 0.9878 0.9836 0.9824 0.0083 0.9788 0.0004 0.0141 0.006
+Table 2: FED scores
 DiaSynth is a synthetic dialogue generation framework designed to address the scarcity of high-quality, large-scale, domain-specific dialogue datasets.
-The framework functions with an LLM as the backbone and CoT as the reasoning mechanism, to simulate diverse real world dialogues that capture the complexity and nuances of human conversations.
-DiaSynth operates by taking in a list of topics from the user based on which the conversations should be generated.
+DiaSynth uses an LLM and CoT reasoning to simulate diverse, nuanced dialogues.
+DiaSynth takes a list of user-provided topics to generate dialogues.
 The users can optionally provide few-shot examples of the format in which they want the dialogue to be generated.
-Dialogues that are generated directly from the given list of topics will be too superficial since the topics given by the user will not be specific.
+Directly generating dialogues from user topics would be too superficial due to their lack of specificity.
 To overcome this lack of specificity, we generate $m$ sub topics for each of the $n$ topics given by the user.
@@
 This is because every dialogue is influenced implicitly by the personas of the people involved in the dialogue and, other characteristics such as the location, emotion and more.
-To increase both variety and depth in the dialogues generated, we generate $p$ personas for each of the $m$ subtopics.
-The dialogues are then generated for all possible combinations of personas belonging to each of the subtopics.
+To enhance variety and depth, we generate $p$ personas per subtopic and create dialogues for all persona-subtopic combinations.
 To further ground the dialogues in various settings and characteristics, we employ CoT reasoning during the generation process.
-DiaSynth employs CoT to reason about the settings and characteristics of a dialogue, as listed in Table 1, ensuring that the dialogues are contextually rich and realistic.
-This three-stage pipeline not only guarantees the quality of the generated dialogues but also allows for exponential scalability, as illustrated programmatically in Appendix A.
+DiaSynth employs CoT to reason about the settings and characteristics of a dialogue, which are listed in Appendix C, ensuring that the dialogues are contextually rich and realistic.
+This three-stage pipeline not only guarantees the quality of the generated dialogues but also allows for exponential scalability, as illustrated programmatically in Appendix B.
 3.1 Subtopic Generation
@@
 Each dialogue will be between two personas who have relevant expertise or perspectives on the given subtopic, allowing the conversation to explore nuances that would otherwise be missed in a generic dialogue setting.
-Notably, during our experiments, the similarity checks indicated no significant overlap among the dialogues, suggesting that the inherent diversity of the dialogues stems from DiaSynth’s multi-stage generation pipeline.
-Figure 1: Overview of the Dialogue Synthesis Framework: The diagram shows the step-by-step process of generating dialogues from user-provided topics, involving sub-topic generation, persona generation, dialogue generation, and filtering based on calculated metrics for downstream tasks.
 3.3 Dialogue Generation
@@
 This step uses an LLM as the backbone and CoT as the reasoning mechanism, allowing the model to simulate dialogues that incorporate various aspects of human interaction.
-In this stage, for each subtopic, DiaSynth generates dialogues by pairing every possible combination of personas within that subtopic.
-The generation process also takes into account the predefined dialogue characteristics as listed in Table 1—such as emotional state, formality level, and familiarity between speakers—to guide the conversation flow and style.
-These characteristics are explicitly defined in the CoT prompt, allowing the LLM to generate dialogues that are both contextually appropriate and realistic.
-Characteristic Description
-Age and Gender Defines the demographic details of both personas, influencing the style, tone, and cultural references in the dialogue.
-Familiarity Level Indicates how well the speakers know each other, which affects the formality, trust, and depth of the conversation.
-Emotional States Specifies the emotional condition of the speakers (e.g., angry, happy, sad), impacting the tone and progression of the dialogue.
-Formality Level Determines the level of politeness or casualness in the dialogue, which depends on context and familiarity between speakers.
-Duration of the Conversation Suggests the intended length of the conversation, which can influence the dialogue’s depth and complexity.
-Communication Medium Defines the medium of the conversation (e.g., face-to-face, phone, text), which affects the dialogue style and expressions used.
-Topic of the Conversation Provides the main subject around which the dialogue is centered, guiding the content and direction of the conversation.
-Location of the Conversation Indicates where the dialogue takes place, adding context that can influence the dialogue’s formality and content.
-Agreement or Disagreement Shows whether the speakers agree or disagree on the topic, which drives the dialogue dynamics and interaction style.
-Natural Dialogue Features Incorporates elements like fillers (e.g., "umm," "uh"), pauses, and slang, adding authenticity and naturalness to the dialogue.
-Table 1: Characteristics of the Dialogue for CoT Prompt
+DiaSynth generates dialogues by pairing all persona-subtopic combinations.
+The process also integrates predefined characteristics (Table 11) like emotional state, formality, and familiarity to guide the flow and style.
+These characteristics are defined in the CoT prompt, guiding the LLM to generate realistic, contextually appropriate dialogues.
+The importance of CoT and the lack of it affects the quality of the dialogues, which is shown quantitatively in Appendix E.
 4 Experimental Setup
-In this section, we detail the experimental setup used to evaluate the effectiveness of DiaSynth.
-Our evaluation focuses on two criteria - quality of the dialogues generated and usability of the dialogues generated for a downstream task.
-Quality of the dialogues is evaluated using metrics such as FED, GPTScore, and G-Eval.
-The usability of the dialogues generated for a downstream task is evaluated by choosing summarization as the downstream task.
-4.1 Quality of the dialogues
-To evaluate the quality of the dialogues, we employ the metrics that have been developed for evaluating the quality of text generated by LLMs. The metrics are as follows,
-  • FED Mehri and Eskenazi (2020) - FED evaluates the quality of dialogue by appending a positive and a negative utterance to the end of the dialogue and measuring the difference in probabilities assigned to the positive and negative utterances.
-  • GPTScore Fu et al.
-(2024) - GPTScore assesses dialogue quality based on various characteristics such as coherence, diversity, and others.
-The evaluation involves asking an LLM whether a dialogue meets a specific criterion.
-The score for each criterion is determined by the probability of the LLM responding affirmatively (’Yes’).
-  • G-Eval Liu et al.
-(2023) - G-Eval evaluates dialogue quality by prompting an LLM to rate the dialogue on a scale of 1 to 3 across multiple criteria.
-The final score is calculated using a weighted average of each rating, with weights derived from the probability distribution of the LLM’s assigned ratings.
-To validate the framework across models and also domains, we generate data using three open source LLMs, one closed source LLM and also use few shot examples from two different dialogue datasets.
-The four LLMs used are - Phi-3 (phi-3-mini-128k-instruct (3.8B)  [^1] Phi-3 HF link) Abdin et al.
-(2024), InternLM-2.5 (internlm2_5-7b-chat (7.74B) [^2] InternLM-2.5 HF link) Cai et al.
-(2024), LLaMA-3 (llama3-8b-instruct (8.03B) [^3] LLaMA-3 HF link) Dubey et al.
-(2024) and GPT-4o.
-The two different dialogue datasets that were used are DialogueSum Chen et al.
-(2021) and SAMSum Gliwa et al.
-(2019)
-4.2 Downstream Task - Summarization
-To evaluate the usability of the dialogues generated by DiaSynth, we choose summarization as the downstream task.
-Summarization is an essential application of dialogue systems, where the goal is to generate concise, and contextually relevant summaries.
-We use three established evaluation metrics—QAGS, BERTScore, and ROUGE-L—to assess the performance of summarization models fine-tuned on DiaSynth-generated data.
-  • QAGS Wang et al.
-(2020) - QAGS (Question Answering Guided Summarization) evaluates the factual consistency of generated summaries by automatically generating questions based on the summary and then comparing the answers obtained from the summary to those from the source dialogue.
-In our work, instead of generating questions from the summaries, we generate them from the source dialogues, allowing us to directly compare the consistency and accuracy of summaries generated from different datasets containing the same dialogues.
-  • BERTScore Zhang* et al. (2020) - BERTScore measures the semantic similarity between the generated summary and the reference summary using contextual embeddings.
-  • ROUGE-L Lin (2004) - ROUGE-L is a widely used evaluation metric that measures the overlap of longest common subsequences (LCS) between the generated summary and the reference summary.
-We fine-tune pretrained models that perform well on summarization such as DistilBART, BART Lewis et al.
-(2020), T5 Raffel et al.
-(2020) and LED Beltagy et al.
-(2020), on DiaSynth-generated dialogues and evaluate their performance using the above metrics.
-We evaluate the effectiveness of DiaSynth in two key aspects: first, by assessing the performance improvement of models fine-tuned on DiaSynth-generated data compared to the pretrained models; and second, by measuring the extent to which DiaSynth-generated data reflects real-world data distribution by comparing the performance of models fine-tuned on DiaSynth data versus those fine-tuned on in-domain data.
-5 Results
-This section discusses the results of the data generated using DiaSynth (quality of the data and usability in downstream tasks) with different LLMs as the backbone and varying few-shot examples.
-Specifically, we utilized Phi-3, InternLM-2.5, LLaMA-3 and GPT-4o as the LLM backbones, and the few-shot examples were sourced from DialogueSum and SAMSum datasets.
-These combinations allow us to evaluate the robustness and adaptability of DiaSynth across different models and few shot examples.
-In total, eight distinct datasets were generated using DiaSynth by pairing each LLM backbone with the two sets of few-shot examples, resulting in all possible combinations.
-For each dataset, DiaSynth was provided with the same 16 broad topics and tasked with generating 6 subtopics for each topic, followed by creating 6 personas for each subtopic.
-The statistics of the datasets generated using DiaSynth, including the number of dialogues, average number of turns, and average number of tokens per turn, are summarized in Table 2.
-LLM Few-Shot Number of Samples Avg. number of turns Avg. number of tokens per turn Diversity (rouge-l)
-Phi-3 DialogueSum 1215 9.13 20.38 0.27
-Phi-3 SAMSum 1410 13.98 13.94 0.27
-InternLM-2.5 DialogueSum 1035 9.23 27.98 0.30
-InternLM-2.5 SAMSum 1135 13.96 19.07 0.29
-LLaMA-3 DialogueSum 1154 6.86 31.99 0.29
-LLaMA-3 SAMSum 1195 10.54 20.41 0.29
-GPT-4o DialogSum 1375 15.16 15.96 0.29
-GPT-4o SAMSum 1380 15.43 13.53 to 0.28
-Table 2: Data Statistics
-5.1 Quality of the Dialogues
-The quality of the synthetic datasets produced by DiaSynth was evaluated using FED, GPTScore, and G-Eval metrics, as detailed in Tables 3, 4, and 5.
-The results reveal distinct variations in performance across different model and dataset configurations, reflecting the unique characteristics of each.
-The FED scores in Table 3 indicate that LLaMA-3 and GPT-4o consistently outperform Phi-3 and InternLM-2.5, particularly in coherence, consistency, and diversity, with LLaMA-3 leading in the DialogSum dataset (e.g., coherence 0.9684, consistency 0.9570, diversity 0.9596) and GPT-4o excelling in the SAMSum dataset, especially in coherence (0.9939) and consistency (0.9878).
-Table 4 further supports LLaMA-3’s adaptability and error-handling abilities, as it outperforms others in diversity (0.0430) and inquisitiveness (0.0415) in DialogSum, while maintaining an edge in SAMSum for diversity (0.0655) and inquisitiveness (0.0622).
-Though GPT-4o remains strong in coherence and consistency, Phi-3 shows competitive coherence in SAMSum (0.0325).
-Finally, Table 5 highlights GPT-4o’s dominance in engagingness and naturalness with perfect scores (3.0) for DialogSum, while InternLM-2.5 stands out in coherence (2.9990) and groundedness (2.9973) for DialogSum, and coherence (2.9983) and groundedness (2.9952) for SAMSum, suggesting it maintains high factual accuracy.
-The contrasting performance of GPT-4o on the DialogSum and SAMSum datasets can be attributed to the differing structures of the dialogues in these datasets.
-DialogSum consists of more formal and structured dialogues, which aligns with the typical response style of GPT-4o, leading to its stronger performance.
-In contrast, SAMSum contains more casual, human-like conversations, which might explain GPT-4o’s relatively poorer performance, as it may not adapt as well to the informal, spontaneous nature of such dialogues.
-Overall, while GPT-4o excels in natural and engaging dialogue, LLaMA-3 offers the most versatility, and InternLM-2.5 provides a strong alternative with high coherence and groundedness.
-The observed superiority of LLaMA-3 over GPT-4o is surprising because an 8 billion 8-bit quantized model not only competes with but also performs better than GPT-4o in certain metrics.
-We hypothesize that this could be due to the way GPT-4o was trained, which might make it more constrained in its responses, whereas LLaMA-3, being an open-source model, operates with fewer restrictions.
-This allows LLaMA-3 to exhibit greater flexibility, diversity, and adaptability in generating dialogues, potentially explaining its better performance in certain metrics.
-These characteristics can be seen in criteria like ’inquisitiveness’ and ’likeability’ in Table 4 and, ’depth’ and ’diverse’ in Table 3.
-These results suggest that for building human-like data generation frameworks, open-source LLMs are a more suitable choice than closed-source LLMs.
-The minimal constraints on response formatting during the training of open-source models enable them to generate more diverse, flexible, and human-like dialogues, making them better suited for tasks requiring natural and conversational interactions.
-coherent error recovery consistent diverse depth likeable understand flexible informative inquisitive
-DIALOGUESUM
-Phi-3 0.9536 0.9440 0.9540 0.9534 0.9521 -0.0005 0.9353 -3.96E-05 0.0009 -0.0033
-InternLM-2.5 0.8439 0.8313 0.8359 0.8353 0.8352 0.0048 0.8278 -0.0046 0.0042 0.0069
-LLaMA-3 0.9684 0.9522 0.9570 0.9596 0.9592 0.0032 0.9453 -0.0063 0.0063 0.0105
-GPT-4o 0.9525 0.9407 0.9417 0.9423 0.9425 0.0121 0.9368 -0.0027 0.0085 0.0144
-SAMSUM
-Phi-3 0.9161 0.9088 0.9199 0.9161 0.9130 -0.0004 0.9014 -0.0024 0.0034 -0.0040
-InternLM-2.5 0.8746 0.8647 0.8734 0.8655 0.8661 0.0033 0.8582 -0.0019 0.0106 0.0028
-LLaMA-3 0.9829 0.9677 0.9757 0.9712 0.9731 0.0003 0.9593 -0.0048 0.0100 0.0029
-GPT-4o 0.9939 0.9876 0.9878 0.9836 0.9824 0.0083 0.9788 0.0004 0.0141 0.006
-Table 3: FED scores
 coherence diversity flexibility understandability inquisitiveness consistency informativeness likeability depth error recovery
@@
 Table 3: GPTScore
+In this section, we detail the experimental setup used to evaluate the effectiveness of DiaSynth.
+Our evaluation focuses on two criteria - quality of the dialogues generated and usability of the dialogues generated for a downstream task.
+Quality of the dialogues is evaluated using metrics such as FED, GPTScore, and G-Eval.
+We evaluate the usability of DiaSynth-generated dialogues by using summarization as the downstream task.
+4.1 Quality of the dialogues
+To evaluate the quality of the dialogues, we employ the metrics that have been developed for evaluating the quality of text generated by LLMs. We use the following metrics:
+  • FED Mehri and Eskenazi (2020) - FED evaluates dialogue quality by comparing the probabilities of appended positive and negative utterances.
+  • GPTScore Fu et al. (2024) - GPTScore assesses dialogue quality by asking an LLM to evaluate criteria like coherence and diversity, with scores based on the probability of affirmative responses.
+  • G-Eval Liu et al. (2023) - G-Eval rates dialogue on a 1-3 scale across criteria, with final scores as a weighted average from the LLM’s probability distribution.
+To validate the framework across models and also domains, we generate data using three open source LLMs, one closed source LLM and also use few shot examples from two different dialogue datasets.
+The open sourced LLMs are - Phi-3 Abdin et al.
+(2024), InternLM-2.5 Cai et al.
+(2024), LLaMA-3 Dubey et al.
+(2024) and the closed source LLM used is GPT-4o.
+The 8-bit quantized versions of the open source LLMs were used for faster experimentation and generation.
+The two different dialogue datasets that were used as few-shot examples are DialogSum Chen et al.
+(2021) and SAMSum Gliwa et al.
+(2019)
+4.2 Downstream Task - Summarization
+To evaluate the usability of the dialogues generated by DiaSynth, we choose summarization as the downstream task.
+Summarization, a key application of dialogue systems, aims to generate concise, contextually relevant summaries.
+We use three established evaluation metrics—QAGS, BERTScore, and ROUGE-L—to assess the performance of summarization models fine-tuned on DiaSynth-generated data.
+  • QAGS Wang et al. (2020) - QAGS evaluates factual consistency by generating questions from the summary and comparing answers to those from the source dialogue.
+  • BERTScore Zhang* et al. (2020) - BERTScore measures semantic similarity between generated and reference summaries using contextual embeddings.
+  • ROUGE-L Lin (2004) - ROUGE-L measures longest common subsequence (LCS) overlap between generated and reference summaries.
+We fine-tune pretrained summarization models like DistilBART, BART Lewis et al.
+(2020), T5 Raffel et al.
+(2020) and LED Beltagy et al.
+(2020), on DiaSynth-generated dialogues and evaluate their performance using the above metrics.
+We evaluate the usability of DiaSynth in two key aspects: first, by assessing the performance improvement of models fine-tuned on DiaSynth-generated data compared to the pretrained models; and second, by measuring the extent to which DiaSynth-generated data reflects real-world data distribution by comparing the performance of models fine-tuned on DiaSynth data versus those fine-tuned on in-domain data.
+5 Results
+This section discusses the results of the data generated using DiaSynth (quality of the data and usability in downstream tasks) with different LLMs and varying few-shot examples.
+Specifically, we utilized Phi-3, InternLM-2.5, LLaMA-3 and GPT-4o as the LLM backbones, and the few-shot examples were sourced from DialogueSum and SAMSum datasets.
+These combinations allow us to evaluate the robustness and adaptability of DiaSynth across different models and few shot examples.
+In total, eight distinct datasets were generated using DiaSynth by pairing each LLM with the two sets of few-shot examples, resulting in all possible combinations.
+For each combination, DiaSynth was provided with the same 16 broad topics and tasked with generating 6 subtopics for each topic, followed by creating 6 personas for each subtopic.
+The statistics of the datasets generated using DiaSynth, including the number of dialogues, average number of turns, and average number of tokens per turn, are summarized in Table 1.
+All the experiments were run on a single A100 GPU with the generation time ranging from 2 hours to 4 hours.
+5.1 Quality of the Dialogues
+The quality of the synthetic datasets produced by DiaSynth was evaluated using FED, GPTScore, and G-Eval metrics, as detailed in Tables 2, 3, and 4.
+The results reveal distinct variations in performance across different model and dataset configurations, reflecting the unique characteristics of each.
+5.1.1 Metric Scores
+FED: The FED scores in Table 2 show that LLaMA-3 and GPT-4o achieve almost a perfect score (+1) in most of the criteria, while Phi-3 and InternLM-2.5 also have decent performances.
+GPT-4o has a clear advantage when it comes to generating likeable dialogues while there is not much separation on other criteria.
+GPTScore: Results illustrated in 3 are surprising in that GPT-4o is the worst performing model on GPTScore, which might require further research while LLaMA-3 clearly dominates the other models.
+G-Eval: Table 4 highlights GPT-4o’s dominance in engagingness and naturalness with perfect scores (3.0) for DialogSum, while InternLM-2.5 stands out in coherence (2.9990) and groundedness (2.9973) for DialogSum, and coherence (2.9983) and groundedness (2.9952) for SAMSum, suggesting it maintains high factual accuracy.
+Dataset-Specific Performance.
+The contrasting performance of GPT-4o on the DialogSum and SAMSum datasets in Table 2 can be attributed to the differing structures of the dialogues in these datasets.
+DialogSum consists of more formal and structured dialogues, which aligns with the typical response style of GPT-4o, leading to its stronger performance.
+In contrast, SAMSum contains more casual, human-like conversations, which might explain GPT-4o’s relatively poorer performance, as it may not adapt as well to the informal, spontaneous nature of such dialogues.
+Overall, while GPT-4o excels in natural and engaging dialogue, LLaMA-3 offers the most versatility, and InternLM-2.5 provides a strong alternative with high coherence and groundedness.
 engagingness naturalness coherence groundedness
@@
 GPT-4o 2.9994 2.9977 2.9982 2.9944
-Table 5: G_EVAL
-5.2 Fine-tuning and Performance Results
-To validate the usability of the synthetic data generated using DiaSynth, we fine-tuned and evaluated several pretrained language models on the task of dialogue summarization.
-The summaries for dialogues generated by different LLMs were created using the corresponding LLMs through prompting.
-The pretrained models used for evaluation include DistilBART, BART, T5, and LED.
-The experimental setup is designed as follows:
-  • All reported metrics are based on the validation and test sets of DialogSum and SAMSum to maintain consistency in evaluation.
-  • To assess the effectiveness of DiaSynth-generated data, we compare the performance of models fine-tuned on DiaSynth data with their corresponding base models, which are evaluated without any fine-tuning.
-  • Given that the training sets of DialogSum and SAMSum are larger than the synthetic datasets generated in this study, we randomly sampled instances from the in-domain training sets to match the size of the DiaSynth-generated data.
-This ensures a fair comparison in terms of training scale.
-  • The primary goal of the experiment is to quantify the percentage improvement offered by models fine-tuned on DiaSynth-generated data compared to their base versions and to evaluate the extent to which the generated data capture the distributional characteristics of the in-domain data.
-The results presented in Table 6 highlight the performance of each model before and after fine-tuning.
-Notably, base models such as DistilBART and LED exhibit relatively high QAGS scores before fine-tuning, primarily because these models tend to repetitively generate content similar to the input dialogues, thereby artificially inflating the factual consistency metric.
-This repetition, while beneficial for preserving factual alignment, does not necessarily translate into improved quality or informativeness of the summaries.
-After fine-tuning on DiaSynth-generated data, models show a significant increase in BERTScore and ROUGE-L compared to their base versions, demonstrating DiaSynth’s effectiveness in enhancing model performance for downstream tasks.
-This result suggests that DiaSynth is valuable for generating dialogue data in domains with limited training resources.
-Notably, models fine-tuned on GPT-4o-generated data performed best for DialogSum, while those fine-tuned on LLaMA-3-generated data excelled on SAMSum, further reinforcing that open-source LLMs are more adept at generating informal, casual dialogues like those in SAMSum.
-The percentage improvement of models fine-tuned on LLaMA-3-generated data was calculated using Equation 1, showing an average improvement of 16.47% across the 24 reported results.
-We chose to calculate on LLaMA-3 because across both quality and usability it performs better than GPT-4o.
-Percentage Improvement=Score After Fine-tuningScore Before Fine-tuningScore Before $\text{Percentage Improvement}=\frac{\text{Score After Fine-tuning}-\text{ScoreBefore Fine-tuning}}{\text{Score Before Fine-tuning}}\times 100$ (1)
-Coverage Percentage=Finetuned Score on DiaSynth DataFinetuned Score on In-Domain Base $\text{Coverage Percentage}=\frac{\text{Finetuned Score on DiaSynth Data}}{\text{Finetuned Score on In-Domain Base Data}}\times 100$ (2)
-To assess the percentage coverage of the distributional characteristics of the in-domain data by the synthetically generated data, we use Equation 2 and the scores of models fine-tuned on LLaMA-3 generated data.
-Across the 24 reported results, the overall coverage percentage of the LLaMA-3 generated data is 90.48%.
-Notably, the QAGS scores of models fine-tuned on synthetic data surpass those of models trained on in-domain data, suggesting that synthetic data can match or even exceed in-domain data performance in some aspects.
-Excluding QAGS, the coverage percentage is calculated to be 77.07%, highlighting that even with moderate LLMs of small scale (3B - 8B), high-quality synthetic dialogue datasets can be effectively created.
-Model Before Finetuning Finetuning on Base Data
+Table 4: G-EVAL
+5.1.2 Strong performance of LLaMA-3
+Model Before Finetuning Finetuning on In-Domain Data
 QAGS BERTScore ROUGE-L QAGS BERTScore ROUGE-L
@@
 Table 5: Performance of models before and after finetuning on in-domain data
+The observed superiority of LLaMA-3 over GPT-4o is surprising because an 8 billion 8-bit quantized model not only competes with but also performs better than GPT-4o in certain metrics.
+We hypothesize that this could be due to the way GPT-4o was trained, which might make it more constrained in its responses, whereas LLaMA-3, being an open-source model, operates with fewer restrictions.
+This allows LLaMA-3 to exhibit greater flexibility, diversity, and adaptability in generating dialogues, potentially explaining its better performance in certain metrics.
+These characteristics can be seen in criteria like ’inquisitiveness’ and ’likeability’ in Table 3 and, ’depth’ and ’diverse’ in Table 2.
+These results suggest that for building human-like data generation frameworks, open-source LLMs are a more suitable choice than closed-source LLMs.
+The minimal constraints on response formatting during the training of open-source models enable them to generate more diverse, flexible, and human-like dialogues, making them better suited for tasks requiring natural and conversational interactions.
+5.2 Fine-tuning and Performance Results
 Model Phi-3 InternLM-2.5 LLaMA-3 GPT-4o
@@
 led-base-16384 0.5883 0.5477 0.2289 0.5457 0.5615 0.2167 0.5917 0.5785 0.2390 0.5738 0.569 0.2298
-Table 7: Performance metrics
-6 Limitations
-Despite the promising results, our approach has some limitations.
-Firstly, different LLMs exhibit varied performance based on the dialogue structure, with certain models like LLaMA-3 performing better for more informal dialogues (e.g., SAMSum) and others like GPT-4o excelling in structured, formal dialogues (e.g., DialogSum).
-This indicates that there is no single model that can universally handle all types of dialogue structures, requiring a tailored choice of LLM depending on the target application.
-Secondly, the generation process may suffer from a lack of knowledge on certain topics, especially in cases where the LLMs were not sufficiently trained on those domains.
-Additionally, our framework relies on zero-shot generation for personas and sub-topics, which, while flexible, can sometimes result in less coherent or less accurate persona simulation, as it is not fine-tuned for specific contexts.
-These limitations suggest directions for future work, such as combining LLMs to leverage their strengths or incorporating more topic-specific training to enhance knowledge coverage.
+Table 6: Performance after finetuning on synthetic data
+To validate the usability of the synthetic data generated using DiaSynth, we fine-tuned and evaluated several pretrained language models on the task of dialogue summarization.
+The summaries for dialogues generated by different LLMs were created using the corresponding LLMs through prompting.
+The pretrained models used for evaluation include DistilBART, BART, T5, and LED.
+The experimental setup is designed as follows:
+  • Metrics are reported on the validation and test sets of DialogSum and SAMSum.
+  • To evaluate DiaSynth-generated data, we compared models fine-tuned on DiaSynth data with their base versions (no fine-tuning).
+  • In-domain training sets were randomly sampled to match the size of the DiaSynth-generated data, enabling fair comparison.
+  • The experiment aimed to quantify performance improvement of DiaSynth-fine-tuned models and assess alignment with in-domain data distributions.
+  • Models were fine-tuned for 2 epochs with a learning rate of 5e-5 and a warmup of 50 steps.
+The results presented in Tables 5 and 6 present the performance of the base models, models finetuned on in-domain data and models finetuned on DiaSynth generated data.
+Models finetuned on DiaSynth data generally improves the performances from the BERTScore and ROUGE-L metrics.
+Surprisingly, for some models (LED and BART) the QAGS scores were higher than the models finetuned on DiaSynth.
+On further exploration, we found out that these models extracted multiple sentences from the given dialogue instead of generating a summary which led to high QAGS scores.
+Comparing models finetuned on in-domain data to those finetuned on DiaSynth data reveals that DiaSynth finetuning generally enhances factual accuracy, with BERTScore and ROUGE-L scores remaining comparable.
+The disparity in BERTScore and ROUGE-L results may be due to format variations.
+Models fine-tuned on in-domain data were evaluated on summaries that matched the training format closely, while DiaSynth-fine-tuned models were trained on LLM-generated summaries and evaluated on human-generated summaries, leading to minor format mismatches.
+Comparison between the different LLMs from Table 6, shows that GPT-4o is better at generating dialogues and summaries that are formal in nature while LLaMA-3 and open source LLMs would be better for generating dialogues that are informal and casual in nature.
+% Improvement $=\frac{\text{After - Before}}{\text{Before}}$ (1a)
+% Coverage $=\frac{\text{Score DiaSynth}}{\text{Score In-domain}}$ (1b)
+Model % Improvement % Covered
+distilbart-cnn 10.96 88.81
+bart-base 9.21 90.6
+t5-base 7.59 93.67
+led-base-16384 2.14 89.68
+Table 7: Summarization results on DialogSum
+Model % Improvement % Covered
+distilbart-cnn 6.07 87.25
+bart-base 16.12 94.35
+t5-base 30.04 87.36
+led-base-16384 15.25 90.91
+Table 8: Summarization results on SAMSum
+To assess the percentage improvement and percentage coverage of the distributional characteristics of the in-domain data by the synthetically generated data, we use Equations 1a and 1b respectively.
+We use the scores of models finetuned on LLaMA-3 generated data because of its dominance in both quality and usability.
+Across the 24 reported results, the overall coverage percentage of the LLaMA-3 generated data is 90.48%.
+Notably, the QAGS scores of models fine-tuned on synthetic data surpass those of models trained on in-domain data, suggesting that synthetic data can match or even exceed in-domain data performance in some aspects.
+Excluding QAGS, the coverage percentage is calculated to be 77.07%.
+In addition to the average percentages, we also present the model wise percentage improvement and coverage in Table 8 and 8.
+The results presented are with respect to the dialogues generated using LLaMA-3 and they illustrate clear improvements for every model, highlighting that even with moderate LLMs of small scale (3B - 8B), high-quality synthetic dialogue datasets can be effectively created across different domains and different dialogue formats.
 6 Conclusion
@@
 Despite certain limitations, such as varying LLM performance across dialogue types and knowledge gaps in zero-shot generation, DiaSynth presents a promising approach to dialogue data generation and offers a valuable resource for future advancements in building more sophisticated and adaptable dialogue systems.
-Appendix A Scalability of DiaSynth
+7 Limitations
+Despite the promising results, our approach has some limitations.
+Firstly, different LLMs exhibit varied performance based on the dialogue structure, with certain models like LLaMA-3 performing better for more informal dialogues (e.g., SAMSum) and others like GPT-4o excelling in structured, formal dialogues (e.g., DialogSum).
+This indicates that there is no single model that can universally handle all types of dialogue structures, but a single SOTA model can give stable and decent results like LLaMA-3.
+Secondly, the generation process may suffer from a lack of knowledge on certain topics, especially in cases where the LLMs were not sufficiently trained on those domains.
+Additionally, our framework relies on zero-shot generation for personas and sub-topics, which, while flexible, can sometimes result in less coherent or less accurate persona simulation, as it is not fine-tuned for specific contexts.
+Since LLMs power DiaSynth, hallucinations and compute-need are two inherent limitations.
+We present a detailed hallucination study in Appendix A, which indicates that though the hallucinations rates are acceptable, there is still scope for improvements.
+These limitations suggest directions for future work, such as combining LLMs to leverage their strengths or incorporating more topic-specific training to enhance knowledge coverage.
+Appendix A Hallucination Study
+In addition to evaluating the quality and usability of dialogues produced by DiaSynth, we conducted a study on the phenomenon of hallucinations within the generated dialogues.
+Hallucinations in language models refer to instances where the output contains misleading or incorrect information or situations where the model repeats the same content.
+To evaluate the occurrence of hallucinations, we compared the generated dialogues with their respective summaries and assessed them using two well-known hallucination benchmarks: SelfCheckGPT Manakul et al.
+[2023] and ChainPoll Friel and Sanyal [2023].
+This analysis provides insights into the prevalence of hallucinations and informs strategies for improving dialogue quality in future iterations of DiaSynth.
+The results are presented in Tables 10 and 10.
+A.1 SelfCheckGPT
+SelfCheckGPT quantifies the self-consistency of LLM outputs by examining agreement across multiple outputs from the same prompt.
+This assessment reveals potential inaccuracies through metrics like SelfCheck-BertScore.
+The SelfCheck-BERTScore results for various models show that hallucination levels are at worst around 25%, which is acceptable but still indicates areas for improvement.
+Across both datasets, Phi-3 demonstrates the most robustness, likely due to its pretraining on structured, textbook-like data, which may contribute to greater consistency and factual accuracy.
+A.2 ChainPoll
+ChainPoll utilizes a chain-of-thought prompting approach to identify hallucinations by iteratively polling the model with structured reasoning prompts.
+This method systematically detects both open-domain and closed-domain hallucinations, where lower scores indicate fewer hallucinations.
+The ChainPoll scores indicate that hallucination levels on these models are generally low, with the best performance seen by GPT-4o on SAMSum, which achieves the lowest score of 0.120, suggesting minimal hallucinations.
+On the other hand, LLaMA-3 scores higher at 0.237 on SAMSum, indicating more frequent hallucinations.
+These findings highlight different models’ strengths in generating accurate and reliable dialogues.
+A.3 Implications for DiaSynth
+The results from both SelfCheckGPT and ChainPoll evaluations suggest that DiaSynth, when leveraging models like Llama 3, is capable of generating dialogues with relatively low hallucination rates.
+However, specific models show variability in performance across datasets, indicating that further enhancements, such as fine-tuning or incorporating additional guardrails, could improve DiaSynth’s robustness in generating reliable dialogues across diverse domains.
+LLM ChainPoll SCGPT-BERTScore
+Phi-3 0.198 0.791
+InternLM-2.5 0.199 0.726
+LLaMA-3 0.205 0.793
+GPT-4o 0.178 0.765
+Table 9: Hallucination calculation for DialogSum few-shot data
+LLM ChainPoll SCGPT-BERTScore
+Phi-3 0.154 0.785
+InternLM-2.5 0.159 0.716
+LLaMA-3 0.237 0.733
+GPT-4o 0.120 0.742
+Table 10: Hallucination calculation for SAMSum few-shot data
+Appendix B Scalability of DiaSynth
 This section illustrates the scalability of DiaSynth with a python program and examples.
@@
 Listing 1: Calculation of Total Dialogs Generated
-A.1 Example 1
+B.1 Example 1
   • Number of topics ($n$): 10
@@
 This setup generates 150 dialogues.
-A.2 Example 2
+B.2 Example 2
   • Number of topics ($n$): 20
@@
 This setup generates 800 dialogues.
-A.3 Example 3
+B.3 Example 3
   • Number of topics ($n$): 15
@@
 This setup generates 4050 dialogues.
-A.4 Scaling Observations
+B.4 Scaling Observations
   • Linear Scaling with Topics and Subtopics: Increasing the number of topics or subtopics results in a linear increase in the total number of dialogues, making it straightforward to expand the scope of dialogue generation.
@@
 These examples illustrate DiaSynth’s potential for rapid and scalable generation of dialogues, which can be tailored to different domains by simply adjusting the input parameters.
+Appendix C Characteristics for the conversation
+Table 11 shows different characteristics that we let the LLMs reason and decide using CoT.
+Before generating the dialogues, the LLMs are prompted to first reason about the various characteristics list for the dialogue given the topic and the personas and how the LLMs reason are illustrated in Appendix D.
+Characteristic Description
+Age and Gender Defines demographic details, influencing style and tone.
+Familiarity Level Affects formality and depth based on relationship between speakers.
+Emotional States Impacts tone and flow based on emotions (e.g., happy, sad).
+Formality Level Determines level of politeness or casualness.
+Duration of the Conversation Suggests the intended length and complexity of dialogue.
+Communication Medium Defines the medium (e.g., face-to-face, phone), influencing style.
+Topic of the Conversation Guides the content and direction of the dialogue.
+Location of the Conversation Adds context influencing formality and content.
+Agreement or Disagreement Drives dialogue dynamics based on agreement level.
+Natural Dialogue Features Adds authenticity with fillers, pauses, and slang.
+Table 11: Characteristics of the Dialogue for CoT Prompt
+Appendix D Example CoT environments
+These examples illustrate how CoT sets the various dialogue characteristics defined in Appendix C.
+As can be seen in the examples, the characteristics for the conversation are completely changed based on the personas and the topics, resulting in more grounded conversation generation.
+Future works can further explore how CoT can be used to further break down to generate even more realistic dialogues.
+(a) CoT Example 1
+(b) CoT Example 2
+Figure 2: Illustration of Example CoT environments
+Appendix E Ablation study on the use of CoT
+In this section of the appendix, we present the impact of CoT in the dialogues generated by DiaSynth.
+We generate two datasets without CoT, using Phi-3 using DialogSum and SAMSum as few-shot examples with 8 topics.
+Tables 13, 13, 15 and 15 compare the scores of dialogues generated with and without CoT for the FED score and GPTScore and it can be clearly that CoT tends to increase the quality of the dialogues generated by DiaSynth.
+We hypothesize that this improvement in quality is due to allowing the LLM to set diverse characteristics for the dialogue before generating the dialogue.
+This illustrates that either manually setting the relevant context or letting the LLM on its own to set the relevant context, we get better outputs, as adding relevant context lowers the probabilities of sequences that are not useful.
+The lower FED scores of CoT generated dialogues in Table 15, might be because of the CoT generated dialogues being longer in length but it needs further research.
+Criteria Without CoT With CoT
+Coherent 0.9507 0.9521
+Error Recovery 0.938 0.9424
+Consistent 0.9424 0.9523
+Diverse 0.9431 0.952
+Depth 0.9451 0.9506
+Likeable 0.0088 -0.0003
+Understand 0.9317 0.9338
+Flexible -0.0013 0.0001
+Informative 0.0032 0.0009
+Inquisitive 0.0095 -0.003
+Table 12: FED scores of dialogues generated with and without CoT for DialogSum few-shot
+Criteria Without CoT With CoT
+Coherence 0.0052 0.0284
+Diversity 0.0120 0.0303
+Flexibility 0.0074 0.0215
+Understandability 0.0056 0.019
+Inquistiveness 0.0162 0.0362
+Consistency 0.0091 0.0366
+Informativeness 0.0099 0.0168
+Likeability 0.0029 0.0209
+Depth 0.0062 0.0115
+Error Recovery 0.0142 0.0242
+Table 13: GPTScore of dialogues generated with and without CoT for DialogSum few-shot
+Criteria Without CoT With CoT
+Coherent 0.9667 0.9125
+Error Recovery 0.9577 0.9051
+Consistent 0.9621 0.9163
+Diverse 0.9586 0.9124
+Depth 0.9589 0.9094
+Likeable 0.0059 -0.0003
+Understand 0.9503 0.8978
+Flexible -0.0022 -0.0023
+Informative 0.0110 0.0035
+Inquisitive 0.0030 -0.0037
+Table 14: FED scores of dialogues generated with and without CoT for SAMSum few-shot
+Criteria Without CoT With CoT
+Coherence 0.0076 0.0324
+Diversity 0.0172 0.0372
+Flexibility 0.0116 0.0259
+Understandability 0.0104 0.0272
+Inquistiveness 0.0201 0.0389
+Consistency 0.0149 0.0415
+Informativeness 0.0148 0.0200
+Likeability 0.0041 0.023
+Depth 0.0078 0.0124
+Error Recovery 0.0163 0.0289
+Table 15: GPTScore of dialogues generated with and without CoT for SAMSum few-shot

### v3

@@
 These results validate DiaSynth’s potential as a robust alternative to traditional data collection methods.
-We open source the code and data generated for future research.
+[^1] https://github.com/ntuspeechlab/DiaSynth We open source the code and data generated for future research.
+DiaSynth: Synthetic Dialogue Generation Framework for Low Resource Dialogue Applications
+Sathya Krishnan Suresh1, Wu Mengjun1, Tushar Pranav2, Eng Siong Chng1 1Nanyang Technological University 2Singapore Institute of Technology Correspondence: sathyakr001@e.ntu.edu.sg
 1 Introduction
@@
 Each dialogue will be between two personas who have relevant expertise or perspectives on the given subtopic, allowing the conversation to explore nuances that would otherwise be missed in a generic dialogue setting.
+We present the impacts of sub-topics and personas in Appendix F
 3.3 Dialogue Generation
@@
 We evaluate the usability of DiaSynth in two key aspects: first, by assessing the performance improvement of models fine-tuned on DiaSynth-generated data compared to the pretrained models; and second, by measuring the extent to which DiaSynth-generated data reflects real-world data distribution by comparing the performance of models fine-tuned on DiaSynth data versus those fine-tuned on in-domain data.
+We also present the results of finetuning BART and T5 on synthetic data and in-domain data on response generation in Appendix H.
 5 Results
@@
 All the experiments were run on a single A100 GPU with the generation time ranging from 2 hours to 4 hours.
+Additional methodolgical details are presented in Appendix G.
 5.1 Quality of the Dialogues
@@
 To evaluate the occurrence of hallucinations, we compared the generated dialogues with their respective summaries and assessed them using two well-known hallucination benchmarks: SelfCheckGPT Manakul et al.
-[2023] and ChainPoll Friel and Sanyal [2023].
+(2023) and ChainPoll Friel and Sanyal (2023).
 This analysis provides insights into the prevalence of hallucinations and informs strategies for improving dialogue quality in future iterations of DiaSynth.
@@
 This section illustrates the scalability of DiaSynth with a python program and examples.
-1def calculate_total_dialogs(n, m, p):
-2 """
-3 Calculate the total number of dialogs generated.
-4
-5 Parameters:
-6 n (int): Number of topics
-7 m (int): Number of subtopics per topic
-8 p (int): Number of personas per subtopic
-9
-10 Returns:
-11 int: Total number of dialogs generated
-12 """
-13 # Calculate the total number of subtopics
-14 total_subtopics = n * m
-15
-16 # Calculate the number of dialogs for each subtopic using combinations of personas
-17 dialogs_per_subtopic = (p * (p - 1)) // 2
-18
-19 # Total dialogs generated across all subtopics
-20 total_dialogs = total_subtopics * dialogs_per_subtopic
-21
-22 return total_dialogs
-Listing 1: Calculation of Total Dialogs Generated
+Algorithm 1 Calculation of Total Dialogs Generated
+1:procedure CalculateTotalDialogs(n, m, p)
+[…]
+[…]
+[…]
+5:     return total_dialogs
+6:end procedure
 B.1 Example 1
@@
 Appendix D Example CoT environments
+Figure 2: CoT Example 1
+Figure 3: CoT Example 2
 These examples illustrate how CoT sets the various dialogue characteristics defined in Appendix C.
@@
 Future works can further explore how CoT can be used to further break down to generate even more realistic dialogues.
-(a) CoT Example 1
-(b) CoT Example 2
-Figure 2: Illustration of Example CoT environments
 Appendix E Ablation study on the use of CoT
@@
 We generate two datasets without CoT, using Phi-3 using DialogSum and SAMSum as few-shot examples with 8 topics.
-Tables 13, 13, 15 and 15 compare the scores of dialogues generated with and without CoT for the FED score and GPTScore and it can be clearly that CoT tends to increase the quality of the dialogues generated by DiaSynth.
+Tables 12, 13, 14 and 15 compare the scores of dialogues generated with and without CoT for the FED score and GPTScore and it can be clearly that CoT tends to increase the quality of the dialogues generated by DiaSynth.
 We hypothesize that this improvement in quality is due to allowing the LLM to set diverse characteristics for the dialogue before generating the dialogue.
@@
 Table 15: GPTScore of dialogues generated with and without CoT for SAMSum few-shot
+Appendix F Ablation Studies on Subtopics and Personas
+To further validate the effectiveness of the DiaSynth framework, we conducted ablation studies by evaluating the impact of removing sub-topics and personas from the data generation pipeline.
+The goal was to assess their contribution to the quality of the generated dialogues.
+For these experiments, we generated approximately 960-1000 dialogues using Phi-3 and compared three settings:
+  • subtopics: Removing sub-topics while keeping personas.
+  • personas: Removing personas while keeping sub-topics.
+  • diasynth: The full DiaSynth-generated data with both personas and sub-topics.
+Metric subtopics personas diasynth
+Coherent 0.9252 0.9584 0.9536
+Error Recovery 0.9022 0.9414 0.944
+Consistent 0.9095 0.9512 0.954
+Diverse 0.9139 0.9512 0.9534
+Depth 0.9193 0.9533 0.9521
+Likeable 0.0069 0.007 -0.0005
+Understandable 0.8918 0.9339 0.9353
+Flexible -0.0038 -0.0042 0
+Informative 0.0096 0.0072 0.009
+Inquisitive 0.0129 0.0063 -0.0033
+Table 16: FED scores with DialogSum as base
+Metric subtopics personas diasynth
+Coherence 0.0118 0.0098 0.0286
+Diversity 0.0246 0.0198 0.031
+Flexibility 0.0239 0.0121 0.0193
+Understandable 0.0137 0.0121 0.0363
+Inquisitive 0.0358 0.0308 0.0363
+Consistent 0.0175 0.0132 0.0369
+Informative 0.0242 0.0155 0.0201
+Likeability 0.0083 0.0076 0.0213
+Depth 0.0216 0.0181 0.0117
+Error Recovery 0.0337 0.0249 0.0342
+Table 17: GPTScore with DialogSum as base
+Metric subtopics personas diasynth
+Coherence 0.0106 0.0109 0.0325
+Diversity 0.0224 0.0191 0.0372
+Flexibility 0.0235 0.0141 0.0395
+Understandable 0.0159 0.0252 0.0415
+Inquisitive 0.0334 0.0137 0.0415
+Consistent 0.0192 0.0415 0.037
+Informative 0.0228 0.0155 0.021
+Likeability 0.0128 0.0075 0.0232
+Depth 0.0187 0.0158 0.0126
+Error Recovery 0.0323 0.0213 0.029
+Table 18: FED scores with SAMSum as base
+Metric subtopics personas diasynth
+Coherence 0.0106 0.0109 0.0325
+Diversity 0.0224 0.0191 0.0372
+Flexibility 0.0235 0.0141 0.0395
+Understandable 0.0159 0.0252 0.0415
+Inquisitive 0.0334 0.0137 0.0415
+Consistent 0.0192 0.0415 0.037
+Informative 0.0228 0.0155 0.021
+Likeability 0.0128 0.0075 0.0232
+Depth 0.0187 0.0158 0.0126
+Error Recovery 0.0323 0.0213 0.029
+Table 19: GPTScore with SAMSum as base
+The ablation studies presented in Tables 16, 17, 18 and 19 demonstrate that the inclusion of both personas and sub-topics significantly enhances the quality of generated dialogues across FED and GPTScore metrics.
+For both DialogSum and SAMSum few-shot examples, dialogues generated with the full DiaSynth framework, incorporating both personas and sub-topics, achieved the highest scores in coherence, diversity, and consistency.
+This indicates that structured dialogue generation benefits from incorporating diverse sub-topics while maintaining persona-driven consistency.
+An interesting observation arises in Table 18, where for the SAMSum dataset, the best-performing configuration involved using only sub-topics without personas.
+This deviation can likely be attributed to the more informal nature of SAMSum dialogues, where structured personas introduce a formal communication style that does not align well with the dataset.
+In contrast, in more structured datasets like DialogSum, the addition of personas provides clear improvements, ensuring dialogue coherence and natural flow.
+Moreover, the largest improvements in quality are seen in coherence, error recovery, and understandability, particularly when both sub-topics and personas are included.
+While sub-topics alone contribute significantly to improving diversity and depth, their combination with personas enhances overall dialogue quality.
+This suggests that dataset characteristics play a crucial role in determining the effectiveness of persona modeling, highlighting the need for adaptive strategies in synthetic dialogue generation.
+Ultimately, these findings reinforce that DiaSynth-generated dialogues are robust and adaptable, providing high-quality synthetic data across both structured and informal conversational settings.
+Appendix G Additional Methodological Details
+G.1 Prompts for Dialogue and Summarization Generation
+The prompts used for dialogue generation were carefully designed to guide the model in producing contextually rich and persona-driven conversations.
+The structure ensures that dialogues exhibit natural human-like characteristics, considering aspects such as speaker familiarity, emotional states, and formality levels.
+For dialogue generation, the following detailed prompt was used:
+You are an expert dialog generator. The following are examples of real-life dialogues.
+example 1
+"$dialogue"-"{dialogue_{1}}"$
+example 2
+"$dialogue"-"{dialogue_{2}}"$
+Use the examples as references and generate a dialogue between people with the following personas:
+$persona1-"{persona_{1}}"$
+$persona2-"{persona_{2}}"$
+Characteristics of the dialogue - to be understood before generating the dialogue: characteristics
+- Assume the values for the characteristics and provide an explanation for choosing those values.
+- These characteristics should be well understood as they implicitly affect and guide the conversation.
+Chain of Thought Reasoning:
+- Before generating the dialogue, reason about the values for each characteristic listed above.
+- Explain in detail how each characteristic will influence a hypothetical dialogue between persons with those characteristics and personas.
+- Ensure that the reasoning considers the interactions between different characteristics (e.g., how familiarity and emotional state might interact).
+- This reasoning and explanation must be included between <cot> and </cot> tags. Do not skip this step.
+- The dialogue generated should be based on the explanation provided.
+This prompt ensures that the model generates structured and realistic dialogues by incorporating a reasoning step before dialogue generation.
+The inclusion of CoT reasoning forces the model to explicitly consider multiple conversational attributes, leading to more coherent and contextually appropriate responses.
+This structured prompting approach enables the generation of high-quality dialogues. The complete list of prompts, including variations for different tasks, is available in our public code repository.
+G.2 Seed Topic Selection
+The seed topics were selected to ensure diversity across conversational scenarios. The topics used in our experiments are as follows:
+Remote Work, Book Recommendations, Fitness Routines, Travel Destinations, Career Development, Movie Reviews, Video Games, Pets, Stock Market, Fashion Trends, Online Education, Mental Wellness, Climate Change, Sports, Artificial Intelligence, Food.
+These topics were chosen from a mix of real-world discussion trends and common conversational themes, ensuring broad coverage and relevance for synthetic dialogue generation.
+G.3 Hyperparameters for Generation
+The hyperparameter settings for different stages of generation are reported below:
+Subtopic Generation:
+Temperature: 0.1 and max tokens: 2048
+Persona Generation:
+Temperature: 0.1 and max tokens: 2048
+Dialogue Generation:
+Temperature: 0.2 and max tokens: 4096
+Appendix H Additional Downstream Task: Response Generation
+Model Before Fine-tuning In-domain Data Llama3 GPT-4o
+t5-base 0.4003 0.6870 0.6572 0.6612
+bart-base 0.5681 0.6875 0.6721 0.6630
+Table 20: BERTScore evaluation of response generation models fine-tuned on synthetic and in-domain datasets.
+To further evaluate the utility of DiaSynth-generated data, we conducted an additional downstream task: response generation.
+This task was included to validate the effectiveness of our synthetic data beyond summarization.
+For these experiments, we selected Llama3 and GPT-4o as base datasets since they demonstrated superior performance across both quality and summarization metrics.
+We employed BERTScore as our evaluation metric due to its effectiveness in measuring the similarity between generated and reference responses.
+The results are presented in Table 20.
+Both models show a significant improvement in BERTScore after fine-tuning on DiaSynth-generated data compared to the pre-trained baseline, highlighting its effectiveness in enhancing model performance.
+Additionally, models fine-tuned on DiaSynth data achieve scores that are close to those fine-tuned on in-domain data.
+For instance, t5-base achieves a BERTScore of 0.6572 on Llama3-generated data and 0.6612 on GPT-4o-generated data, compared to 0.6870 for in-domain fine-tuning.
+These results indicate that DiaSynth-generated data serves as a viable alternative for fine-tuning response generation models, performing comparably to in-domain data, even in low-resource scenarios.